# Exploring the Practicality of Generative Retrieval on Dynamic Corpora

### v4

@@ 3.1 DynamicIR Task Setup
To conduct these experiments, we assume that we have an initial corpus $C_{\textit{initial}}$ and a newly introduced corpus $C_{\textit{new}}$, and datasets of query-document pairs $R_{\textit{initial}}$ and $R^{\prime}_{\textit{new}}$ from $C_{\textit{initial}}$ and $C_{\textit{new}}$, respectively. Unlike $R_{\textit{initial}}$, $R^{\prime}_{\textit{new}}$ consists of pseudo-queries, which are generated from $C_{\textit{new}}$ using docT5qeury (detailed explanation is in Section 3.2). These query-document pairs are used for supervised learning. Moreover, we assess the retrieval performance with two types of evaluation sets, $Q_{\textit{initial}}$ and $Q_{\textit{new}}$, where the answers to the questions are within $C_{\textit{initial}}$ and $C_{\textit{new}}$, respectively. Some questions between $Q_{\textit{initial}}$ and $Q_{\textit{new}}$ are identical except for the timestamps, which necessitates the retrieval of different passages. Each set is employed to assess the forgetting of initial knowledge McCloskey and Cohen (1989); Kirkpatrick et al. (2017) and the acquisition of new knowledge.
@@ 3.2 Benchmark
To evaluate the performance of retrieval models in a dynamic scenario, we employ StreamingQA Liška et al. (2022) designed for temporal knowledge updates. Unlike other benchmarks on temporal retrieval Deveaud et al. (2023), StreamingQA is the only benchmark that includes both the timestamps of question asked time and document publication dates, which is critical for considering the temporal dynamics. The temporal information is prepended to the text in the format of ‘Today is Wednesday, May 6, 2020. [question]’ for question, and ‘Thursday, February 7, 2019. [document text]’ for documents Liška et al. (2022). The dataset covers 14 years and includes over 50 million passages, more than double the content size of Wikipedia used in DPR (21M).
@@
The original StreamingQA dataset lacks query-document pairs from ${C_{\textit{new}}}$, making it challenging to explore training-based updates including supervised learning. To address this, we generate additional 90,000 queries from ${C_{\textit{new}}}$. To make this, we employ a trained model similar to the one used in docT5111https://github.com/castorini/docTTTTTquery for query generation. The size of this additional dataset $R^{\prime}_{\textit{new}}$ is similar to that of $R_{\textit{initial}}$. Examples of generated dataset are in Table A.7 and details of the query construction are explained in Appendix A.5.
@@ 4.1 Retrieval Models
Although our main focus are on transformer-based semantic search models to explore corpora adaptation, we also evaluate BM25 (Lucene) [^2] https://github.com/castorini/pyserini, a traditional lexical matching retriever.
@@ 4.2 Effective Continual Pretraining of GR
When continually pretraining GR on $C_{\textit{new}}$, we employ LoRA Hu et al. (2021) widely recognized for its training efficiency. To better target key parameters for incorporating new knowledge, we analyze which parameters undergo the most significant change during the acquisition of new knowledge. This analysis is inspired by works on model merging Ansell et al. (2023); Wortsman et al. (2022). We refer these crucial parameters in learning new knowledge as Dynamic Parameters (DPs).
@@ 5 Results and Analysis
In this section, we showcase the adaptability and robustness of GR and DE and provide an analysis on effective continual pretraining approach and utilizing $R^{\prime}_{\textit{new}}$ during the training-based update. We also discuss the performance of BM25 in dynamic environments.
@@ 6 Computation & Memory Efficiency
We analyze the inference FLOPs FLOPs (Floating Point Operations) is the number of floating-point arithmetic calculations. of DE and GR to assess their computational efficiency. The FLOPs function and detailed calculations are in Appendix A.7. As shown in Table 3, our results reveal that GR requires 2 times fewer computations per instance over DE. Figure 4 illustrates that GR offers superior efficiency as the number of instances increases. Moreover, unlike DE, which exhibits $\mathcal{O}(N)$ complexity, where $N$ represents the corpus size, GR maintains a constant $\mathcal{O}(1)$ complexity.
@@
There is a difference in the concept of indexing between DE and GR. For DE, this involves embedding, which converts the corpus into representations using an encoder. In GR, indexing refers the data processing of document identifiers to constrain beam search decoding, ensuring the generation of valid identifiers. Note that we process data without applying sharding.
As shown in Table 3, our results exhibit that GR (3.1h) requires 6$\times$ less time than DE (20.4h) for indexing $C_{\textit{initial}}$ and $C_{\textit{new}}$. The crucial aspect of indexing is that unlike GR, DE necessitates re-indexing the entire corpus each time whenever the model is updated, irrespective of the corpus update.
@@
Inference process can be divided into two stages: (1) loading a pre-indexed corpus and (2) retrieving, which includes query embedding and search. We classify the former as offline latency ($T_{\textit{offline}}$) and the latter is referred to as online latency ($T_{\textit{online}}$). We measure both. $T_{\textit{online}}$ in 3 is reported for a single instance.
Table 3 shows GR is 10 times faster than DE when retrieving from updated corpora for $T_{\textit{offline}}$. Unlike DE, which stores each passage representation in vector form, GR does not need much time to load the index since it stores knowledge within its parameters. For $T_{\textit{online}}$, however, GR is 20 times slower than DE using faiss-gpu. Although DE requires 2 times more inference flops, it seems that the FAISS Johnson et al. (2019) module contributes significantly to the inference speed of DE.
@@
Table 3 indicates that GR has 4 less storage requirements over DE for updated corpora. Notably, the memory requirements for DE are directly affected by the corpus size, as they store representations of all documents in vector form outside the retrieval model. In contrast, GR has minimal dependence on the corpus size by storing knowledge in its internal parameters.
@@ 7 Conclusion
In this work, we explore the practicality of GR in terms of adaptability to new corpora, robustness, and inference efficiency, compared to DE. By establishing a DynamicIR setup, we showcase how retrieval models perform in dynamic environments where knowledge evolves over time with and without parameter updates. Our findings show GR’s superior performance in these three aspects and highlight its potential as a practical IR system.

### v5

@@ 3.1 DynamicIR Task Setup
To conduct these experiments, we assume that we have an initial corpus $C_{\textit{initial}}$ and a newly introduced corpus $C_{\textit{new}}$, and datasets of query-document pairs $R_{\textit{initial}}$ and $R^{\prime}_{\textit{new}}$ from $C_{\textit{initial}}$ and $C_{\textit{new}}$, respectively. Unlike $R_{\textit{initial}}$, $R^{\prime}_{\textit{new}}$ consists of pseudo-queries, which are generated from $C_{\textit{new}}$ using docT5query (detailed explanation is in Section 3.2). These query-document pairs are used for supervised learning. Moreover, we assess the retrieval performance with two types of evaluation sets, $Q_{\textit{initial}}$ and $Q_{\textit{new}}$, where the answers to the questions are within $C_{\textit{initial}}$ and $C_{\textit{new}}$, respectively. Some questions between $Q_{\textit{initial}}$ and $Q_{\textit{new}}$ are identical except for the timestamps, which necessitates the retrieval of different passages. Each set is employed to assess the forgetting of initial knowledge McCloskey and Cohen (1989); Kirkpatrick et al. (2017) and the acquisition of new knowledge.
@@ 3.2 Benchmark
To evaluate the performance of retrieval models in a dynamic scenario, we employ StreamingQA Liška et al. (2022) designed for temporal knowledge updates. Unlike other benchmarks on temporal retrieval Deveaud et al. (2023), StreamingQA is the only benchmark that includes both the timestamps of question-asked time and document publication dates, which is critical for considering the temporal dynamics. The temporal information is prepended to the text in the format of ‘Today is Wednesday, May 6, 2020. [question]’ for question, and ‘Thursday, February 7, 2019. [document text]’ for documents Liška et al. (2022). The dataset covers 14 years and includes over 50 million passages, more than double the content size of Wikipedia used in DPR (21M).
@@
The original StreamingQA dataset lacks query-document pairs from ${C_{\textit{new}}}$, making it challenging to explore training-based updates including supervised learning. To address this, we generate additional 90,000 queries from ${C_{\textit{new}}}$. To make this, we employ a trained model similar to the one used in docT5query [^1] https://github.com/castorini/docTTTTTquery for query generation. The size of this additional dataset $R^{\prime}_{\textit{new}}$ is similar to that of $R_{\textit{initial}}$. Examples of generated dataset are in Table A.7 and details of the query construction are explained in Appendix A.5.
@@ 4.1 Retrieval Models
Although our main focus is on transformer-based semantic search models to explore corpora adaptation, we also evaluate BM25 (Lucene) [^2] https://github.com/castorini/pyserini, a traditional lexical matching retriever.
@@ 4.2 Effective Continual Pretraining of GR
When continually pretraining GR on $C_{\textit{new}}$, we employ LoRA Hu et al. (2021) widely recognized for its training efficiency. To better target key parameters for incorporating new knowledge, we analyze which parameters undergo the most significant change during the acquisition of new knowledge. This analysis is inspired by works on model merging Ansell et al. (2023); Wortsman et al. (2022). We refer to these crucial parameters in learning new knowledge as Dynamic Parameters (DPs).
@@ 5 Results and Analysis
In this section, we showcase the adaptability and temporal robustness of GR and DE and provide an analysis on the effective continual pretraining approach and utilizing $R^{\prime}_{\textit{new}}$ during the training-based updates. We also discuss the performance of BM25 in dynamic environments.
@@ 6 Computation & Memory Efficiency
We analyze the inference FLOPs FLOPs (Floating Point Operations) is the number of floating-point arithmetic calculations. of GR and DE to assess their computational efficiency. The FLOPs function and detailed calculations are in Appendix A.7. As shown in Table 3, our results reveal that GR requires 2 times fewer computations per instance over DE. Figure 4 illustrates that GR offers superior efficiency as the number of instances increases. Moreover, unlike DE, which exhibits $\mathcal{O}(N)$ complexity, where $N$ represents the corpus size, GR maintains a constant $\mathcal{O}(1)$ complexity.
@@
There is a difference in the concept of indexing between GR and DE. For DE, this involves embedding, which converts the corpus into representations using an encoder. In GR, indexing refers the data processing of document identifiers to constrain beam search decoding, ensuring the generation of valid identifiers. Note that we process data without applying sharding.
As shown in Table 3, our results exhibit that GR (3.1h) requires 6 times less time than DE (20.4h) for indexing $C_{\textit{initial}}$ and $C_{\textit{new}}$. The crucial aspect of indexing is that unlike GR, DE necessitates re-indexing the entire corpus each time whenever the model is updated, irrespective of the corpus update.
@@
Inference process can be divided into two stages: (1) loading a pre-indexed corpus and (2) retrieving, which includes query embedding and search. We classify the former as offline latency ($T_{\textit{offline}}$) and the latter is referred to as online latency ($T_{\textit{online}}$). We measure both. $T_{\textit{online}}$ in Table 3 is reported for a single instance.
Table 3 shows GR is 10 times faster than DE when retrieving from updated corpora for $T_{\textit{offline}}$. Unlike DE, which stores each passage representation in vector form, GR does not need much time to load the index since it stores knowledge within its parameters. For $T_{\textit{online}}$, however, GR is 20 times slower than DE using faiss-gpu. Although DE requires 2 times more inference flops, it seems that the FAISS Johnson et al. (2019) module contributes significantly to the inference speed of DE. For GR, there is potential for further speedup using techniques such as FlashAttention Dao et al. (2022); Dao (2023), speculative decoding Leviathan et al. (2023), and a Mixture of Experts Shazeer et al. (2017); Lepikhin et al. (2020); Kudugunta et al. (2021).
@@
Table 3 indicates that GR has 4 times less storage requirements over DE for updated corpora. Notably, the memory requirements for DE are directly affected by the corpus size, as they store representations of all documents in vector form outside the retrieval model. In contrast, GR has minimal dependence on the corpus size by storing knowledge in its internal parameters.
@@ 7 Conclusion
In this work, we explore the practicality of GR in terms of adaptability to new corpora, temporal robustness, and inference efficiency, compared to DE. By establishing a DynamicIR setup, we showcase how retrieval models perform in dynamic environments where knowledge evolves over time, with and without parameter updates. Our findings indicate that GR better adapts to changing knowledge and less forgets by leveraging the power of language models. In terms of temporal robustness, GR retrieves information effectively regardless of data characteristics, while DE struggles with temporal information. Additionally, GR has fewer inference FLOPs, reduced indexing time, and lower memory requirements. By comprehensively exploring GR across these three aspects, we underscore its potential as an IR system.